SpaceRef

SpaceRef


NOAA GOES SATOPS Morning Report - June 19, 2003 Day 170

Status Report From: National Oceanic and Atmospheric Administration
Posted: Friday, June 20, 2003

DMSP 

On June 18, a restack of PR 3 and 4 was performed on OPS47. 

On June 18, a 10K 2 point dwell was performed on gyro 1 and 3's motor current on OPS49. 

All DMSP operations were nominal over the past 24 hours. 

POES 

All POES operations were nominal over the past 24 hours. 

GOES 

At 1136Z on June 18, GOES-12 telemetry was lost (and the schedule suspended) as a result of a TACTS2 crash at CDA.  The crew failed over to TCS1, reconfigured for command/telemetry and restarted the command schedule at 1142Z.  The GOES-12 1139Z Southern Hemi image was missed as a result. 

At 0325Z on June 19, a GOES-9 frame break was observed and attributed to loss of signal synchronization. 

GOES-10 will need an E/W maneuver (as a result of residual E/W drift from the N/S maneuver on Tuesday), likely next Thursday. 

GOES-8 continues its westward drift at 1.06 deg/day.  The drift table with the estimated current position can be found at http://www.oso.noaa.gov/goes/GOES-8_Drift_Table.txt.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.